• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 04:10:53 on April27,2003

    Entry number 98754

    keyword=sieve slit removed, more BGR study

    N. Liyanage and A. Deur took access to remove the sieve slit on
    the RHRS. Now we do runs with 3-foil C12 target, 1-foil, and no target.
    Run 20306/1306 -- 3 foil target, 2 uA. DAQ saturates.
    20307/1307 - same, 1 uA, higher psfactors.
    Next we put in the 1-foil target.
    20308/1308 -- 1 foil target.
    T1 = 74 kHz, T2 = 530, T3 = 230, T4 = 11
    20309/1309 -- empty target
    After this run, we take access to replace the sieve slit and put
    in the target collimator, with additional collimation to cut the Be window.